OpenStudy (sburchette):

Stokes' Theorem I'm trying to evaluate the line integral around the curve obtained by intersecting the cylinder x^2+z^2=1 with the plane y=3. The vector field is F=<3xz, e^(xz), 2xy> and the curve is oriented counterclockwise when viewed from the right.

OpenStudy (sburchette):

I parameterized the surface needed to use stokes' thm as r(u,v)=<u*cos(v), 3, u*sin(v)> where 0 <=u<=1 and 0<=v<=2*pi. The curl of F and the cross product of the partials of r(u,v) are quite long, but I got an answer of 6*pi which is far from the answer in the solution key.

OpenStudy (amistre64):

F(r) . r' right?

OpenStudy (sburchette):

For the line integral it is, but we are supposed to use Stokes' theorem

OpenStudy (amistre64):

i can never keep the names straight

OpenStudy (amistre64):

what is stokes thrm by chance?

OpenStudy (slaaibak):

Flux is F.n

OpenStudy (slaaibak):

The line integral in space is equal to the flux of the surface

OpenStudy (sburchette):

\[\int\limits_{C} F.dr = \int\limits \int\limits_{S} curl(F).dS\]

OpenStudy (slaaibak):

what was your normal vector? <0,y,0? ?

OpenStudy (amistre64):

oy, its way to late for me to be doing anything remotely sane with that :/

OpenStudy (anonymous):

If the plane intersects the cylinder you know the cross section's shadow on the x-y plane is a circle right? So you need only to integrate around the circle? And since it's a surface integral r if FIXED at r=1 not to be integrated over. So integrate over z and phi (the polar angle).

OpenStudy (anonymous):

I believe.

OpenStudy (anonymous):

Especially with y=3, its not even slanted so it will DEFINITELY only give you a circle of constant radius.

OpenStudy (sburchette):

The formula I have been following states that flux integral over a paramterized surface is \[\int\limits \int\limits_{D} F(r).(r_u \times r_v) dA\] This, in effect, eliminates having to compute the normal directly.

OpenStudy (anonymous):

And since you're only integrating with scalars you don't need to worry about integrating over a "position" vector which would NOT give you a circle if the plain was slanted. And you still compute the normal directly, you just don't normalize it.

OpenStudy (anonymous):

And I guess it would be the x-z plane not x-y.

OpenStudy (sburchette):

Integrating along the curve would likely be more simple, but we are supposed to use Stokes' theorem instead of directly computing the line integral

OpenStudy (slaaibak):

but the normal is quite obvious though, isn't it? It just points in the y-direction. so <0,1,0> is fine

OpenStudy (sburchette):

Yes, I meant we don't compute the nomalized normal, the wording gets tricky ;)

OpenStudy (anonymous):

And I misread, since you want to calculate the line integral you only need to integrate over phi. So:  So: \[\Phi(\phi)=(- \cos(\phi),3,\sin(\phi))\] That is your parameterization for that circle. So \[0 \le \phi \le 2 \pi\] And then you have: \[\oint \vec{F} \cdot d \vec{l}=\int\limits_0^{2 \pi} \vec{F}(\vec{\Phi}(\phi)) \cdot \vec{\Phi}'(\phi)d \phi\]